# Reason3D: Searching and Reasoning 3D Segmentation via Large Language Model

### v1

@@ 4.3 Training 3D LLM
The loss function for Reason3D comprises two essential components: the LLM loss $\mathcal{L}_{llm}$ and the segmentation mask loss $\mathcal{L}_{mask}$. The overall combination is represented as:
$\mathcal{L}=\mathcal{L}_{llm}+\mathcal{L}_{mask}.$ (5)
@@
In particular, The LLM loss, $\mathcal{L}_{llm}$, embodies the linguistic aspects through an auto-regressive cross-entropy loss for text generation, incorporating cross-entropy loss ${\rm CE}$ for each token:
@@
where $\hat{\mathbf{Y}}_{txt}$ represents the ground truth word token. In addition, the mask loss $\mathcal{L}_{mask}$ aims at encouraging the model to generate high-quality segmentation masks. This loss is computed using a binary cross-entropy (BCE) loss and DICE loss for all superpoints, which is represented as:
@@
where $\hat{\mathbf{M}}_{*}$ means the ground truth segmentation mask for region-level and object-level superpoints. For the object-level mask $\hat{\mathbf{M}}_{seg}$, we utilize the mask of the object that we target. For the region-level mask $\hat{\mathbf{M}}_{loc}$, we mark the points as the foreground points if the distance between any point and the object center is smaller than threshold $\tau$.
@@ 5 Experiments
5.1 Experimental Setting
Datasets. Our training data includes three main types of datasets: (1) For the 3D expressive referring segmentation task, we use ScanRefer [7] and Sr3D datasets [1]. (2) For the 3D question answering task, we utilize ScanQA dataset [4]. (3) For the 3D reasoning segmentation task, we use the Reason3D datasets, which we annotated from both the ScanNetV2 and Matterport3D datasets. The results of 3D QA and more details are included in the Appendix.
Model Architecture. We use a pre-trained Sparse 3D U-Net [38] to extract point-wise features. For the language learning model, we employ FlanT5 [12], maintaining most of its pre-trained weights frozen, except for adapting the weights for the newly-added location and segmentation tokens. Our QFormer is constructed by the BLIP-2 [26] architecture, incorporating 1408-dimensional features.
Evaluation Metrics. For the 3D expressive referring segmentation and 3D reasoning segmentation tasks, the primary evaluation metrics are Mean Intersection over Union (mIoU), which quantifies the average overlap between the predicted and true 3D volumes, and Accuracy at k Intersection over Union (Acc@kIoU). This latter metric measures the proportion of descriptions for which the predicted mask overlaps the ground truth with an IoU greater than k, where k is set at thresholds of 0.25 and 0.5, thus assessing the model’s performance at varying levels of precision. For the QA task, the evaluation metrics include BLEU-4 [30], ROUGE-L [27], METEOR [5], and CIDEr [43] to ensure robust answer matching. These metrics evaluate the precision, fluency, and semantic accuracy of the responses.
5.2 3D Reasoning Segmentation Results
We present the results of our 3D reasoning segmentation in Table 1. Our model significantly surpasses previous models in tasks that require advanced reasoning and a deep understanding of world knowledge, achieving a notable increase in mean Intersection over Union (mIoU). Differing from typical 3D referring segmentation tasks, this task demands not just an understanding of spatial relationships but also robust reasoning capabilities and a broad comprehension of context.
Our model excels at interpreting long sentence queries and managing 3D reasoning segmentation tasks, outperforming open-vocabulary segmentation methods such as OpenScene [31] and OpenMask3D [39]. We also compare its performance against a two-stage method, where FlanT5 generates a short text output for the query in the first stage, followed by segmentation using either OpenScene or OpenMask3D. Our method surpasses this approach by leveraging more expressive hidden embeddings, which provide a richer representation than solely relying on text as an intermediary.
Compared with leading 3D referring segmentation models, specifically 3D-STMN [45] fine-tuned on the Reason3D dataset, we observe that although 3D-STMN excels in direct referencing, it struggles with indirect queries. In contrast, our model adeptly navigates these challenges by integrating large language models, thereby demonstrating superior adaptability and performance.
In addition, we compare our model to another two-stage method that combines Llama2 [20] and CLIP [35], both fine-tuned on the Reason3D dataset. This approach begins with Llama2 generating an output vocabulary, which CLIP then processes to extract textual features. These features act as queries interacting with point cloud features via a mask decoder (we exploit the same point features and mask decoder as our method), ultimately producing the segmentation mask results. The results demonstrate that our approach significantly outperforms this two-stage method, which is completely decoupled and relies solely on the textual outputs from the LLM model.
@@ 5.1 Experimental Setting
Table 1: 3D Reasoning Segmentation Results on Reason3D dataset. The evaluation metric is accuracy at IoU 0.25, IoU 0.5 and mIoU.
@@
0.25 0.50 mIoU 0.25 0.50 mIoU 0.25 0.50 mIoU
@@
Method Room Num = 1$\sim$2 Room Num = 3$\sim$4 Room Num >= 5
0.25 0.50 mIoU 0.25 0.50 mIoU 0.25 0.50 mIoU
Reason3D (w/o HMD) 25.23 10.32 15.56 12.84 5.50 8.23 8.26 2.52 5.33
Region Seg + Reason3D (w/o HMD) 26.98 13.21 17.02 19.21 8.21 11.67 11.96 4.78 7.21
Reason3D (w/ HMD) 29.82 16.97 18.81 22.25 11.93 14.12 16.06 7.34 10.35
Table 3: 3D Reasoning Segmentation Results on Matterport3D dataset with different room numbers to validate the effectiveness of the design of the proposed Hierarchical Mask Decoder (HMD). The evaluation metric is IoU@0.25, IoU@0.5 and mIoU.
@@
To demonstrate the effectiveness of our model in the 3D express referring segmentation task, we compare our approach against current state-of-the-art methods on ScanRefer validation set in Table 2. It can be observed that our approach significantly outperforms 3D-STMN [45] in overall performance. Considering the limited focus on the 3D referring segmentation task in the existing literature, we extend our comparison to several 3D grounding approaches that only predict 3D bounding boxes. These methods can predict segmentation masks by simply extracting points within the predicted boxes. Notably, our approach vastly outperforms the LLM-based method, 3D-LLM [17], which struggles with accurately locating 3D boxes and effectively extracting segmentation masks.
@@ 5.4 Ablation Study
Effectiveness of Hierarchical Mask Decoder. We present experimental results in Table 3 to demonstrate the effectiveness of our proposed Hierarchical Mask Decoder (HMD) design in searching objects across varying numbers of rooms. Since the Reason3D dataset primarily focuses on single-room scenarios, we extend it to multi-room settings by reusing a subset of the annotated Matterport3D [6] data to include a varying number of regions. We opt not to utilize ScanNet, as it features only single-room scenes.
We establish three variants of our Reason3D model: (1) Reason3D without the Hierarchical Mask Decoder, (2) Reason3D combined with a region segmentation module that first predicts and segment the target room’s region then applying the Reason3D model for the segmented region, and (3) Reason3D equipped with the Hierarchical Mask Decoder. We can validate the effectiveness of the proposed Hierarchical Mask Decoder (HMD) by comparing the first and last models, especially when the number of rooms exceeds five. The effectiveness of the HMD is likely due to the difficulty of integrating language embeddings with point cloud features as the number of points increases (i.e., in more rooms scenarios). In contrast, the hierarchical strategy of the HMD, which uses an additional token to learn a coarse region as a prior, effectively manages the output segmentation masks. In addition, the second model segments point clouds from the region of interest before using the Reason3D framework without the HMD. This method outperforms the first model but faces optimization challenges within the training pipeline of the two-stage approach.
@@
Effectiveness of Different Designs. Table 4 shows the ablation experiments of different choices. Table 4(a) proves that the superpoints pooling operation is essential for our pipeline since it helps to reduce the training complexity and enables the effective training of the pipeline. Also, the average pooling for the superpoints can achieve better performance. Table 4(b) presents the performance impact of various components of segmentation loss. Using either Binary Cross-Entropy (BCE) loss or Dice loss alone leads to significantly reduced performance. In contrast, combining Dice loss and BCE loss results in the most favorable outcomes. Table 4(c) presents the impact of different numbers of decoder layers. We use 6 layers for the decoder as the default number.
@@ 5.5 Visualization Results
Figure 4 displays the visualization results of our Reason3D model for the 3D reasoning segmentation task, highlighting our model’s proficiency in accurately generating segmentation masks based on the query. Additional visualization results are included in the Appendix.
@@ 6 Conclusion
This study presents Reason3D, an innovative framework that leverages Large Language Models (LLMs) for enhanced scene understanding, capable of generating textual responses and segmentation predictions. To demonstrate the effectiveness of our approach, we introduce the novel task of 3D reasoning segmentation, which requires interpreting implicit human instructions within three-dimensional scenes. We have developed a hierarchical mask decoder that improves the accuracy of final mask predictions by initially identifying a broad region likely to contain the target object, which then serves as a basis for further refinement. Our extensive experiments on the ScanNetV2 and Matterport3D datasets demonstrate outstanding performance in various complex tasks, including 3D reasoning segmentation, 3D referring segmentation, and question answering.
Appendix
Appendix A Reason3D Dataset
@@ A.1 Dataset annotation
Each 3D scene in the Reason3D dataset consists of a textual query and a binary segmentation mask to identify the target objects. As mentioned in the paper, we utilize ScanNetV2 [13] and Matterport3D [6] as our data source. Considering single room space as one scene, we first extract the instance object annotation and the room type information from these two datasets as the tag of 3D scenes. After that, we utilize these tags as parts of the text prompt to incorporate with GPT-4. The illustration of the prompt construction process is shown in Table 5, and some samples utilized for prompting are shown in Table 6.
messages = [ {"role":"system", "content": You are an AI visual assistant, and you are seeing a 3D scene. What you see is provided with several words to represent objects with tag <objects>, describing the scene you are looking at, and also the room type <type> to describe the type of the scene. Design a question <question> that can be answered confidently with the <answer> from one of the provided objects in <objects>. Please do not ask any <question> that cannot be answered confidently. Each question should have one clear answer that is most relevant, without ambiguity or multiple possible answers in the list of description words. Please include complex questions relevant to the scene’s content, such as inquiries into the background knowledge of objects or discussions about events related to these objects. Avoid questions about uncertain or unclear details. The question should be natural.} ] for  sample in few_shot_samples:
@@ A.2 Dataset Statistics
As detailed in the main paper, the Reason3D dataset incorporates the ScanNetV2 and Matterport3D datasets. We adhere to their official training and validation splits for data annotation. Specifically, the Matterport3D dataset provides 934 training samples and 837 validation samples. Meanwhile, the ScanNetV2 dataset contributes 405 training samples and 308 validation samples.
Appendix B Experiments
@@ B.1 Implementation Details
We execute the models on two NVIDIA Tesla A100 GPUs, using a batch size of 32 for training and a batch size of 1 for inference. We employ the AdamW optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and a weight decay of 0.05. Additionally, we apply a linear warm-up strategy to the learning rate during the initial 1,000 steps, increasing it from $10^{-8}$ to $10^{-4}$, followed by the cosine decay schedule. All experiments are implemented with the Pytorch framework.
B.2 Data Statistics
@@
ScanNet [13], a comprehensive 3D indoor dataset, covers diverse environments including apartments and various room types. The dataset is structured into 1201 training scenes, 312 validation scenes, and 100 testing scenes. We adopt the training configuration as SPFormer [38] for training our scene encoder on the ScanNet training dataset.
@@
ScanQA [4] is a dataset for 3D question answering task based on ScanNet, consisting of 25,563 question-answer pairs on 562 scenes for training and 4,675 question-answer pairs on 71 scenes for validation.
B.3 3D Question Answering Results
In addition to excelling in 3D reasoning and referring tasks, our approach also performs well in 3D question answering tasks. We present our results on the ScanQA validation set in Table 8, where we observed a significant improvement in evaluation metrics over both baseline methods and the recent LLM-based method, 3D-LLM. Our approach not only answers questions accurately but also visualizes the related segmentation masks to further demonstrate the effectiveness.
B.4 3D Visual Grounding Results.
Although our primary focus is on 3D segmentation, our method also effectively predicts 3D bounding boxes as supplementary outputs, facilitating comparison with 3D visual grounding methods. To generate a 3D bounding box for a referred object, we first apply DBSCAN [15] to eliminate noisy points, and then calculate the minimum and maximum XYZ coordinates from the points within the segmentation mask to form the 3D box. As demonstrated in Table 8, our approach not only outperforms recent 3D visual grounding methods but also significantly surpasses LLM-based methods, such as 3D-LLM [17], which struggle to integrate textual and numerical data to accurately localize objects in 3D space.

### v2

@@ 4.3 Training Reason3D
The loss function for Reason3D comprises two essential components: the LLM loss $\mathcal{L}_{\rm{llm}}$ and the segmentation mask loss $\mathcal{L}_{\rm{mask}}$. The overall combination is represented as:
$\mathcal{L}=\mathcal{L}_{\rm{llm}}+\mathcal{L}_{\rm{mask}}.$ (5)
@@
In particular, The LLM loss, $\mathcal{L}_{\rm{llm}}$, embodies the linguistic aspects through an auto-regressive cross-entropy loss for text generation, incorporating cross-entropy loss ${\rm CE}$ for each token:
@@
where $\hat{\mathbf{Y}}_{\rm{txt}}$ represents the ground truth word token. In addition, the mask loss $\mathcal{L}_{\rm{mask}}$ aims at encouraging the model to generate high-quality segmentation masks. This loss is computed using a binary cross-entropy (BCE) loss and DICE loss for all superpoints, which is represented as:
@@
where $\hat{\mathbf{M}}_{*}$ means the ground truth segmentation mask for region-level and object-level superpoints. For the object-level mask $\hat{\mathbf{M}}_{\rm{seg}}$, we use the mask corresponding to the specific object we are targeting. For the region-level mask $\hat{\mathbf{M}}_{\rm{loc}}$, we designate the points as foreground points if the distance between any point and the object’s center is smaller than threshold $\tau$ or we select the points within the specific room for hierarchical searching task.
@@ 5 Experiments
Table 1: 3D Reasoning Segmentation Results. The evaluation metric is accuracy at IoU 0.25, IoU 0.5 and mIoU.
Method Room Num = 1$\sim$2 Room Num = 3$\sim$4 Room Num $\geq$ 5
Acc@0.25 Acc@0.50 mIoU Acc@0.25 Acc@0.50 mIoU Acc@0.25 Acc@0.50 mIoU
(a) FlanT5 [12] + OpenScene [35] 17.65 3.95 12.89 11.27 1.02 7.69 6.22 0.97 2.21
(b) Reason3D-base 25.23 10.32 15.56 12.84 5.50 8.23 8.26 2.52 5.33
(c) Region Seg + Reason3D-base 26.98 13.21 17.02 19.21 8.21 11.67 11.96 4.78 7.21
(d) Reason3D 29.82 16.97 18.81 22.25 11.93 14.12 16.06 7.34 10.35
Table 2: 3D Hierarchical Searching Results on Matterport3D dataset with different room numbers. Reason3D-base refers to the full Reason3D model without the proposed [LOC] token and region decoder. The evaluation metric is IoU@0.25, IoU@0.5 and mIoU.
@@
Acc@0.25 Acc@0.50 mIoU Acc@0.25 Acc@0.50 mIoU Acc@0.25 Acc@0.50 mIoU
@@
Ablation Acc@0.25 Acc@0.50
(a) w/o region sup. ($\mathcal{L}_{\rm{mask}_{loc}}$) 14.27 6.33
(b) hard thresholding for $\mathbf{M}_{\rm{loc}}$ 20.35 9.98
(c) Eq. 4 $\rightarrow$ concat. $\mathcal{F}_{\rm{seg}}$ and $\mathcal{H}_{\rm{loc}}(\mathbf{M}_{\rm{loc}})$ 21.53 10.98
(d) Reason3D (full model) 22.25 11.93
Table 4: The effect of different designs in Hierarchical Mask Decoder (HMD) on Matterport dataset. We use room number = 3$\sim$4 as the main experiments.
5.1 Experimental Setting
Datasets. Our training data includes three main types of datasets: (1) For the 3D expressive referring segmentation task, we use ScanRefer [7] and Sr3D datasets [1]. (2) For the 3D question answering task, we utilize ScanQA dataset [4]. (3) For the 3D reasoning segmentation task, we construct the Reason3D dataset from ScanNetV2 and Matterport3D datasets. The results of 3D QA and more details are included in the supplementary materials.
Model Architecture. We use a pre-trained Sparse 3D U-Net [45] to extract point-wise features. For the language learning model, we employ FlanT5 [12], maintaining most of its pre-trained weights frozen, except for adapting the weights for the newly-added location and segmentation tokens. Our Interactor is constructed following BLIP-2 [29], incorporating 1408-dimensional features.
Evaluation Metrics. For the 3D expressive referring segmentation and 3D reasoning segmentation tasks, the primary evaluation metrics are Mean Intersection over Union (mIoU), which quantifies the average overlap between the predicted and true 3D volumes, and Accuracy at k Intersection over Union (Acc@kIoU). This latter metric measures the proportion of descriptions for which the predicted mask overlaps the ground truth with an IoU greater than k, where k is set at thresholds of 0.25 and 0.5, thus assessing the model’s performance at varying levels of precision.
5.2 3D Reasoning Segmentation Results
Table 1 presents the results of 3D reasoning segmentation, where our model significantly outperforms previous methods, achieving a notable increase in mean Intersection over Union (mIoU). Unlike typical 3D referring segmentation tasks, this task demands not only spatial understanding but also robust reasoning and contextual comprehension.
Our model excels at interpreting long sentence queries and managing 3D reasoning segmentation tasks, outperforming open-vocabulary segmentation methods like OpenScene [35] and OpenMask3D [46], which primarily use vocabulary as the query. We also compare it to the two-stage methods, where FlanT5 [12] generates a short vocabulary output followed by segmentation with OpenScene [35] or OpenMask3D [46]. Our approach surpasses these by leveraging more expressive hidden embeddings, offering a richer representation than relying solely on text as an intermediary.
Compared to leading 3D referring segmentation models like 3D-STMN [52] fine-tuned on the Reason3D dataset, we find that while 3D-STMN excels in direct referencing, it struggles with indirect queries. In contrast, our model, with its integration of large language models, shows superior adaptability and performance in these scenarios.
We also compare our model to a two-stage method that combines Llama2 [22] and CLIP [40], both fine-tuned on the Reason3D dataset. In this approach, Llama2 [22] generates a vocabulary output, which CLIP [40] converts into textual features. These features then interact with the same point features and mask decoder used in our Reason3D to produce segmentation masks. The results show that our approach significantly outperforms this two-stage method, which is fully decoupled and relies solely on the textual outputs from the LLM.
5.3 3D Hierarchical Searching Results
For the 3D hierarchical searching task, an extension of 3D reasoning segmentation task, our goal is to segment target objects within a larger space (e.g., multiple rooms) rather than a single room. The task involves specifying a target room where the model must locate the object, such as finding the TV in the bedroom, as shown in Figure 3. Given that the Reason3D dataset mainly focuses on single-room scenarios, we extend it to multi-room settings by reusing a subset of annotated Matterport3D [6] data. We chose not to use ScanNet, as it features only single-room scenes.
Table 2 presents different results: (a) a two-stage baseline that uses an LLM [12] to parse the instruction, and applies an open-world segmentation model [35]; (b) Reason3D-base (without Region Decoder), (c) Reason3D combined with a region segmentation model that first segments the target room’s region, then applies the Reason3D-base for the segmented region (without location prompt), and (d) Reason3D (full model). Reason3D (d) outperforms the two-stage baseline (a) due to its effective design. Comparing (b) and (d) shows that the Hierarchical Mask Decoder (HMD) significantly boosts performance by using an additional token to guide coarse region learning, effectively managing segmentation masks, especially as point cloud complexity increases. Additionally, baseline (c) improves upon (b) but faces optimization challenges in the two-stage training pipeline compared to the full model (d).
@@ 5.1 Experimental Setting
To demonstrate the effectiveness of our model in the 3D express referring segmentation task, we compare Reason3D against state-of-the-art methods on the ScanRefer validation set, as shown in Table 3. Our approach significantly outperforms 3D-STMN [52] in overall performance. Given the limited focus on 3D referring segmentation in the literature, we also compare our model to several 3D grounding approaches that predict only 3D bounding boxes, which can generate segmentation masks by extracting points within the predicted boxes. Notably, our approach vastly outperforms the LLM-based method, 3D-LLM [19], which struggles with accurately locating 3D boxes and effectively extracting segmentation masks.
@@ 5.5 Ablation Study
Effectiveness of different design in HMD. Besides the Reason3D-base (without the proposed coarse-to-fine approach) in Table 2, we also conduct an ablation study on the Matterport3D dataset for the 3D hierarchical searching task to validate the impact of various designs in our Hierarchical Mask Decoder (HMD), as shown in Table 4. When the region supervision term ($\mathcal{L}_{\rm{mask}_{loc}}$) is removed (a), the model’s performance drops significantly, achieving only 14.27% accuracy at a 0.25 IoU, which indicates that region supervision plays a crucial role in guiding the HMD decoder to learn the priors for segmentation predictions. Comparing (b) and (d), we observe that hard thresholding performs worse than probability-based region mask prediction. Hard thresholding discards uncertainty, while using probability as a prior retains valuable information, enabling more nuanced and accurate segmentation decisions. Furthermore, comparing (c) and (d) suggests that summation is more effective than concatenation for combining point features with learned location priors in Equation 4, likely due to its more integrated and simpler feature representation.
@@
Effectiveness of superpoints. Table 5(a) proves that the superpoints pooling operation is essential for our pipeline since it helps to reduce the training complexity and enables the effective training of the pipeline. Also, the average pooling for the superpoints can achieve better performance.
Effectiveness of different segmentation loss. Table 5(b) presents the performance impact of various components of segmentation loss. Using either Binary Cross-Entropy (BCE) loss or Dice loss alone leads to significantly reduced performance. In contrast, combining Dice loss and BCE loss results in the most favorable outcomes.
Effectiveness of different decoder layers. Table 5(c) presents the impact of different numbers of decoder layers. We use six layers for the decoder as the default number.
@@ 5.6 Visualization Results
Figure 4 displays the visualization results of our Reason3D model for the 3D reasoning segmentation task, highlighting our model’s proficiency in accurately generating segmentation masks based on the query. Additional visualization results are included in the supplementary materials.
@@ 6 Conclusion
This paper presents Reason3D, a framework that leverages Large Language Models (LLMs) for enhanced scene understanding, generating textual responses and segmentation predictions. We introduce the novel task of 3D reasoning segmentation, requiring interpreting implicit human instructions within three-dimensional scenes. A hierarchical mask decoder is proposed to enhance mask prediction by first identifying a broad region likely to contain the target object, which then serves as feature priors for further refinement. Extensive experiments on the ScanNetV2 and Matterport3D datasets demonstrate outstanding performance across tasks like 3D reasoning segmentation, 3D hierarchical searching, 3D referring segmentation, and question answering.
Supplementary Material
We provide additional details on our dataset in Section A. Section B elaborates on the implementation and presents further results. Sections C to E offer comprehensive discussions on potential failure cases, limitations, and the broader impact of our method.
A Reason3D Dataset
@@ A.1 Dataset Annotation
Each 3D scene in the Reason3D dataset consists of a textual query and a binary segmentation mask to identify the target objects. As mentioned in the paper, we utilize ScanNetV2 [13, 43] and Matterport3D [6] as our data sources. Considering single room space as one scene, we first extract the instance object annotation and the room type information from these two datasets as the tags of 3D scenes. After that, we utilize these tags as parts of the text prompt to incorporate with GPT-4. The illustration of the prompt construction process is shown in Table 6, and some samples utilized for prompting are shown in Table 7.
messages = [ {"role":"system", "content": You are an AI visual assistant, and you are seeing a 3D scene. What you see is provided with several words to represent objects with tag <objects>, describing the scene you are looking at, and also the room type <type> to describe the type of the scene. Design a question <question> that can be answered confidently with the <answer> from one of the provided objects in <objects>. Please do not ask any <question> that cannot be answered confidently. Each question should have one clear answer that is most relevant, without ambiguity or multiple possible answers in the list of description words. Please include complex questions relevant to the scene’s content, such as inquiries into the background knowledge of objects or discussions about events related to these objects. Avoid questions about uncertain or unclear details. The question should be natural.} ] for  sample in few_shot_samples do
@@ A.2 Dataset Statistics
The collected Reason3D dataset incorporates the ScanNetV2 [13] and Matterport3D [6] datasets. We adhere to their official training and validation splits for data annotation. Specifically, the Matterport3D dataset provides 934 training samples and 837 validation samples. Meanwhile, the ScanNetV2 dataset contributes 405 training samples and 308 validation samples.
A.3 3D Hierarchical Searching Extension
The 3D hierarchical searching task extends the reasoning-based 3D segmentation task by incorporating a specified target room type where the queried object should be located. As detailed in the main paper, we utilize only a subset of the Matterport3D dataset, chosen for its diversity in room types. In this task, we include the template "In <ROOM_TYPE>" to specify the target room, where <ROOM_TYPE> represents the various room categories defined in the Matterport3D dataset. For experiments involving different numbers of rooms, we expand the target room’s space by including its neighboring rooms, adjusting this space according to the specified number of rooms.
B Experiments
@@ B.1 Implementation Details
Our models are primarily executed on two NVIDIA RTX A6000 GPUs, using a batch size of 16 during training and 1 during inference. The AdamW optimizer is employed with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and a weight decay of 0.05. Additionally, we implement a linear warm-up strategy for the learning rate during the initial 1,000 steps, gradually increasing it from $10^{-8}$ to $10^{-4}$, followed by a cosine decay schedule. All experiments are conducted using the PyTorch framework.
B.2 Used Datasets
@@
ScanNet [13, 43], a comprehensive 3D indoor dataset, covers diverse environments including apartments and various room types. The dataset is structured into 1201 training scenes, 312 validation scenes, and 100 testing scenes.
@@
ScanQA [4] is a dataset for the 3D question answering task based on ScanNet, consisting of 25,563 question-answer pairs on 562 scenes for training and 4,675 question-answer pairs on 71 scenes for validation.
B.3 Training
Following the approach in [45, 52], we first train a sparse 3D UNet [17] as our backbone and compute superpoint features using pre-computed superpoints [16, 28]. Once the backbone is fully trained, its weights are frozen. For each target task, we then pre-train the entire network using data from the other tasks, followed by task-specific fine-tuning.
B.4 Instruction Template
In this section, we present the instructions and outputs used for task-specific templates. Following previous works [56, 10], we utilize a “human:“ identifier to initiate the instruction, followed by an “assistant:“ identifier for the LLM-generated response. We use <scene> to represent the token corresponding to the point cloud scene. The tokens <SEG> and <LOC> are used for segmentation and location, respectively, as part of the prompting process for generating segmentation results, as described in the main paper. Below are examples for various tasks, where {description} refers to the target object’s description, {question} represents a query based on the given scene, and {answer} is the corresponding response.